# Efficient Fourier Filtering Network with Contrastive Learning for UAV-based Unaligned Bi-modal Salient Object Detection

### v2

@@ III-B Semantic contrastive alignment loss (SCAL)
In UAV-based unaligned bi-modal images, differences in resolution and fields of view lead to variations in the scale and position of salient objects in the spatial dimension. It is important that both modalities maintain consistent spatial position representation [1]. To address the challenge of aligning heterogeneous modalities, we note that contrastive learning [25, 26, 27, 28, 29] has recently demonstrated strong performance. Contrastive alignment [28, 29] is essentially a process of mutual refinement between modalities, where each modality benefits from knowledge exchange to enhance scene comprehension and achieve alignment. In addition, the high-level features that contain rich semantic information can directly affect the model’s understanding of the scene and the reconstruction of detailed features. With these factors in mind and without adding inference costs, we propose SCAL to achieve semantic-level spatial positional alignment of the two modalities and enhance single-modality representation through contrastive communication.
As illustrated in the diagram of SCAL in Fig. 2, since the receptive field of the thermal modality is small, we align each element in $f_{4}^{t}\in\mathbb{R}^{B\times C\times H\times W}$ as a base with the corresponding element in $f_{4}^{r}\in\mathbb{R}^{B\times C\times H\times W}$, where $B$, $C$, $H$, and $W$ represent batch size, channels, height, and width of features, respectively. First, we compress the number of channels of $f_{4}^{r}$ and $f_{4}^{t}$ to $1$ using a $3\times 3$ depthwise separable convolution [61] layer $DConv_{3}(\cdot)$ and a normalization layer $Norm(\cdot)$:
@@
$l_{u}=\begin{cases}1&\text{if }s_{u}>t\\ 0&\text{otherwise}\end{cases},$ (4)
where $l_{u}$ represents the label of $p_{u}$. Specifically, it is classified as a positive sample when $s_{u}$ is larger than $t$; otherwise, it is classified as negative. To bring positive samples closer to $a$ and push negative samples farther away, we employ InfoNCE [62] to measure the distance between samples:
@@ III-C Synchronized Alignment Fusion (SAF)
Although bi-modal features are aligned at deep layers, they still remain slightly offset at high-resolution shallow layers. In addition, the distribution of features varies significantly between modalities due to their distinct physical properties. Effectively fusing these complementary features is essential for achieving BSOD [1, 18, 12, 52, 46, 22]. To address these challenges, we propose SAF, which uses an FFT-based multiple-filtering strategy to fuse bi-modal complementary features and align them in both spatial and channel dimensions.
As shown in Fig. 2, SAF consists of normalization layers, residual connections, a feed-forward network, and a token mixer. To capture the local relevance, we equip a $DConv_{3}(\cdot)$ in the feed-forward network [21]. For the token mixer, we introduce the efficient group filter (EGF). In EGF, we first integrate the normalized bi-modal features $\tilde{f_{i}^{r}}$ and $\tilde{f_{i}^{t}}$ into the same feature $f_{i}^{a}$, reducing the computational cost for feature alignment and fusion. This process is denoted as:
@@
where $+$ is element-wise addition. After applying a StarReLU [63] layer $SR(\cdot)$, we perform an FFT on the integrated feature $f_{i}^{a}$, then multiply it by the group filter $\mathcal{G}(\cdot)$, where each channel is linearly coupled. The group channel weights in $\mathcal{G}(\cdot)$ are determined by applying global average pooling $GAP(\cdot)$ to $f_{i}^{a}$. The use of channel-weighted $\mathcal{G}(\cdot)$ enhances the sensitivity of capturing salient details and spatial shifts in the spatial dimension, thereby facilitating the fusion and alignment of bi-modal features. Finally, the output $\tilde{f_{i}^{s}}$ of EGF is obtained by inverse FFT (IFFT). The whole process is expressed as:
@@
where $\times$ is element-wise multiplication. The $N$ global filters [58] that make up $\mathcal{G}(\cdot)$ are denoted as $\{G_{1},G_{2},\ldots,G_{N}\}$. The channel-weighted $\mathcal{G}(\cdot)\in\mathbb{C}^{B\times C\times H\times(\frac{W}{2}+1)}$ is defined as:
@@ III-D Decoder
In encoder-decoder structured networks, the distribution of features varies across levels. Low-level features capture basic patterns, such as edges and textures, while high-level features represent more abstract concepts and semantic information [31, 30]. To utilize these diverse representations, we introduce a simple decoder that fits these features gradually from high to low levels. Each decoder layer consists of a $DConv_{3}(\cdot)$, a $Norm(\cdot)$, and ReLU function, denoted as $DNR(\cdot)$. The whole process is represented as:
@@
For the loss function, we employ the commonly used BCE [64] and IoU [65] losses to supervise the learning of $\mathcal{S}$, and employ SCAL to supervise the alignment of $f_{4}^{r}$ and $f_{4}^{t}$, which can be mathematically described as:
@@ IV-A Datasets
To comprehensively validate our proposed AlignSal, we first conducted experiments on the UAV-based unaligned bi-modal salient object detection (BSOD) task. To further assess its generalization, we extended our evaluation to three additional bi-modal tasks: weakly aligned BSOD, aligned BSOD, and remote sensing change detection (RSCD).
1) UAV-based Unaligned BSOD: We conducted extensive experiments on the UAV RGB-T 2400 dataset [1], which contains 2400 instantly captured unaligned RGB-T image pairs falling at different times of the day and different seasons of the year. Each image pair has a corresponding pixel-wise annotation. Following [1], all BSOD models were trained on the UAV RGB-T 2400 training set and tested on the testing set.
2) Weakly Aligned BSOD: We employed unaligned-VT821, unaligned-VT1000, and unaligned-VT5000 datasets [53] for the experiments. These three datasets contain 6821 RGB-T image pairs and were obtained by performing random affine transformations from the three corresponding aligned datasets VT821 [36], VT1000 [66], VT5000 [11]. Following [1, 53], all BSOD models were trained on the unaligned-VT5000 training set and tested on the remaining datasets.
3) Aligned BSOD: We utilized VT821 [36], VT1000 [66], and VT5000 [11] datasets, which collectively contain 6821 manually aligned RGB-T image pairs in the natural scene. Following [12, 22], all BSOD models were trained on the VT5000 training set and tested on the remaining datasets.
4) RSCD: To assess the generalization of AlignSal in high-altitude scenarios, we conducted experiments on the LEVIR-CD+ dataset [67], which comprises 985 remote sensing image pairs documenting significant building changes. Each image has an ultra-high resolution of 1024 $\times$ 1024 pixels at 0.5 meters per pixel. All RSCD models were trained on the dataset’s training set and evaluated on its testing set.
@@ IV-B Evaluation Metrics
To fairly evaluate model performance for BSOD tasks, we used six widely recognized evaluation metrics. E-measure [68] ($Em$) assesses the joint contribution of pixel-level accuracy and image-level matching. S-measure [69] ($Sm$) combines region-aware and object-aware evaluation to quantify the structural similarity. F-measure [70] (${F_{\beta}}$) balances precision and recall to provide a harmonic mean assessment of accuracy. While weighted F-measure [71] (${wF_{\beta}}$) adjusts precision and recall by pixel importance for a refined saliency evaluation. Mean absolute error [72] ($\mathcal{M}$) calculates the average pixel-wise absolute difference between prediction and ground truth. We also used Precision-recall (PR) curve to show the trade-off between precision and recall across various thresholds.
When evaluating model performance for the RSCD task, we employed five additional metrics [73], including recall (Rec.), precision (Pre.), overall accuracy (OA), F1 score (F1), and IoU [65].
@@
TABLE I: Quantitative Comparison of E-measure ($Em$), S-measure ($Sm$), F-measure (${F_{\beta}}$), Weighted F-measure (${wF_{\beta}}$), and Mean Absolute Error ($\mathcal{M}$) Between Our AlignSal and Sixteen State-of-the-Art BSOD Models on the UAV RGB-T 2400 Dataset [1]. Type A, U, and UU Stand for Aligned, Unaligned, and UAV-based Unaligned BSOD Models, Respectively. The Best Results are Labeled Red and the Second Best Results are Labeled Blue. We Tested the FPS of Different Models on an NVIDIA RTX4060 Laptop GPU.
@@
CATNet23 [19] A Swin-B 0.7420 0.8155 0.6331 0.5189 0.0182 262.6M 341.8G 19.7
PICRNet23 [52] A Swin-T 0.8609 0.7685 0.6129 0.6083 0.0191 112.0M 27.1G 30.3
@@
LAFB24 [23] A Res2Net-50 0.9633 0.9040 0.8471 0.8161 0.0068 453.0M 139.7G 29.3
DCNet22 [53] U VGG-16 0.9276 0.8292 0.7500 0.7997 0.0090 24.1M 207.2G 14.5
MROS23 [1] UU Res2Net-50 0.9747 0.8963 0.8506 0.8494 0.0063 90.4M 45.3G 12.2
AlignSal (Ours) UU CDFFormer-S18 0.9654 0.9085 0.8599 0.8277 0.0062 27.1M 22.9G 30.8
@@
$Em\uparrow$ $Sm\uparrow$ $wF_{\beta}\uparrow$ $F_{\beta}\uparrow$ $\mathcal{M}\downarrow$uparrow $Em\uparrow$ $Sm\uparrow$ $wF_{\beta}\uparrow$ $F_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $Em\uparrow$ $Sm\uparrow$ $wF_{\beta}\uparrow$ $F_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $Em\uparrow$ $Sm\uparrow$ $wF_{\beta}\uparrow$ $F_{\beta}\uparrow$ $\mathcal{M}\downarrow$
@@
CATNet23 [19] 0.7734 0.8331 0.6627 0.5645 0.0183 0.7511 0.8227 0.6520 0.5393 0.0166 0.7281 0.7953 0.6079 0.5021 0.0207 0.8031 0.8946 0.7514 0.5975 0.0054
PICRNet23 [52] 0.8678 0.7632 0.6067 0.6233 0.0239 0.8600 0.7671 0.6175 0.6101 0.0185 0.8469 0.7461 0.5842 0.5830 0.0219 0.9206 0.8566 0.7510 0.7240 0.0061
@@
LAFB24 [23] 0.9769 0.9142 0.8759 0.8531 0.0057 0.9673 0.9115 0.8650 0.8332 0.0067 0.9576 0.8914 0.8297 0.7990 0.0081 0.9819 0.9511 0.9199 0.8842 0.0020
DCNet22 [53] 0.9529 0.8369 0.7748 0.8229 0.0090 0.9424 0.8314 0.7639 0.8134 0.0089 0.9185 0.8120 0.7285 0.7839 0.0105 0.9740 0.8886 0.8490 0.8813 0.0038
@@ IV-C Implementation Details
We trained our model on an NVIDIA Tesla P100 GPU, using the Adam optimizer [74] with a batch size of 8. Our model utilized the CDFFormer-S18 [57] as the encoder, which is pre-trained on ImageNet. The initial learning rate was set to $3\times 10^{-5}$ and decreased by a factor of 10 every 100 epochs, over a total of 300 epochs. The input resolution was set to $384\times 384$.
@@ IV-D Model Comparison
We compared our AlignSal with fourteen state-of-the-art aligned BSOD models, one state-of-the-art unaligned BSOD model, and one state-of-the-art UAV-based unaligned BSOD model. The aligned BSOD models include MIDD [13], CSRNet [14], MoADNet [15], SwinNet [16], OSRNet [17], TNet [18], MGAI [12], CATNet [19], PICRNet [52], LSNet [46], HRTNet [20], CAVER [21], WaveNet [22], and LAFB [23]. The unaligned BSOD model is DCNet [53]. The UAV-based unaligned BSOD model is MROS [1]. All saliency maps were obtained from the authors or generated by the open-source code and were evaluated by the Saliency Evaluation Toolbox [^2] https://github.com/jiwei0921/Saliency-Evaluation-Toolbox.
Quantitative Evaluation. To fairly evaluate the performance of our AlignSal on the UAV RGB-T 2400 dataset [1], we showed the comparison of $Em$, $Sm$, $wF_{\beta}$, $F_{\beta}$, $\mathcal{M}$,
Figure 3: Visual comparison of our AlignSal with state-of-the-art BSOD models in several challenges, including fast UAV movement ($1^{st}$, $2^{nd}$, and $6^{th}$ rows) and fast object movement ($8^{th}$ row) in object blurring scenes, low illumination ($1^{st}$ and $8^{th}$ rows), street light exposure ($3^{rd}$ and $9^{th}$ rows), and extreme low illumination ($4^{th}$ row) in illumination changing scenes, small objects ($2^{nd}$, $4^{th}$, $5^{th}$, and $8^{th}$ rows), out-of-view ($1^{st}$, $2^{nd}$, $6^{th}$, and $8^{th}$ rows), multiple objects($1^{st}$-$4^{th}$, $7^{th}$-$9^{th}$ rows), scale variation ($8^{th}$ row), and center bias ($4^{th}$ and $6^{th}$ rows) in object changing scenes, rain ($8^{th}$ row) and snow ($9^{th}$ row) in weather changing scenes. GT represents ground truth.
Figure 4: (a) Precision-recall (PR) and (b) F-measure-threshold (FT) curves of BSOD models on the UAV RGB-T 2400 dataset.
and three complexity metrics (i.e., Params, FLOPs, and FPS) of our AlignSal with the above BSOD models in Table I. Notably, with real-time inference speed and lower complexity, our model achieved superior performance compared to the state-of-the-art BSOD models. Specifically, compared to the existing top-performing UAV-based unaligned model MROS [1], our model surpassed it in three out of five metrics (i.e. an improvement of 1.4%, 1.1%, and 1.6% in $Sm$, $wF_{\beta}$, and $\mathcal{M}$, respectively), while being much lighter (i.e. 70% reduction in Params) and having much faster inference speed (i.e. half the FLOPs and 2.5 times the FPS). Compared to the existing best aligned BSOD model LAFB [23], our model outperformed it by 0.2%, 0.5%, 1.5%, 1.4%, and 8.8% in $Em$, $Sm$, $wF_{\beta}$, $F_{\beta}$, and $\mathcal{M}$, respectively, while having a 94.1% and 83.6% reduction of Params and FLOPs. In addition, we also provided the comparison of PR and F-measure-threshold (FT) curves in Fig. 4. We can observe that the PR curve of our model is closest to (1,1), and the area under the FT curve is the largest as well.
@@
Qualitative Comparison. As shown in Fig. 3, we present a visual comparison of our AlignSal with existing state-of-the-art BSOD models across several challenging scenes, including fast UAV movement ($1^{st}$, $2^{nd}$, and $6^{th}$ rows) and fast object movement ($8^{th}$ row) in object blurring scenes; low illumination ($1^{st}$ and $8^{th}$ rows), street light exposure ($3^{rd}$ and $9^{th}$ rows), and extreme low illumination ($4^{th}$ row) in illumination changing scenes; small objects ($2^{nd}$, $4^{th}$, $5^{th}$, and $8^{th}$ rows), out-of-view ($1^{st}$, $2^{nd}$, $6^{th}$, and $8^{th}$ rows), multiple objects($1^{st}$-$4^{th}$, $7^{th}$-$9^{th}$ rows), scale variation ($8^{th}$ row), and center bias ($4^{th}$ and $6^{th}$ rows) in object changing scenes; rain ($8^{th}$ row) and snow ($9^{th}$ row) in weather changing scenes. As shown, a scene often contains multiple challenges and AlignSal can completely capture objects in these extreme scenes. This is attributed to our proposed SCAL and SAF and the way they cooperate within AlignSal.
TABLE IV: Ablation Studies of the Key Components in AlignSal on the UAV RGB-T 2400 Dataset. The Best Results are Labeled Red.
@@
SCAL aligns and mutually refines bi-modal features at the semantic level using a contrastive strategy, while SAF aligns the two modalities at the pixel level by using the powerful filtering mechanism to effectively explore the bi-modal complementary information and capture fine details.

### v3

@@ III-B Semantic contrastive alignment loss (SCAL)
In AAV-based unaligned bimodal images, differences in resolution and fields of view lead to variations in the scale and position of salient objects in the spatial dimension. It is important that both modalities maintain consistent spatial position representation [10315195]. To address the challenge of aligning heterogeneous modalities, we note that contrastive learning [radford2021learning, yang2021taco, he2023align, jiang2024t, song2024contrastalign] has recently demonstrated strong performance. Contrastive alignment [jiang2024t, song2024contrastalign] is essentially a process of mutual refinement between modalities, where each modality benefits from knowledge exchange to enhance scene comprehension and achieve alignment. In addition, the high-level features that contain rich semantic information can directly affect the model’s understanding of the scene and the reconstruction of detailed features. With these factors in mind and without adding inference costs, we propose SCAL to achieve semantic-level spatial positional alignment of the two modalities and enhance single-modality representation through contrastive communication.
As illustrated in the diagram of SCAL in Fig. 2, since the receptive field of the thermal modality is small, we align each element in $f_{4}^{t}\in\mathbb{R}^{B\times C\times H\times W}$ as a base with the corresponding element in $f_{4}^{r}\in\mathbb{R}^{B\times C\times H\times W}$, where $B$, $C$, $H$, and $W$ represent batch size, channels, height, and width of features, respectively. First, we compress the number of channels of $f_{4}^{r}$ and $f_{4}^{t}$ to $1$ using a $3\times 3$ depthwise separable convolution [chollet2017xception] layer $DConv_{3}(\cdot)$ and a normalization layer $Norm(\cdot)$:
@@
[…]
where $l_{u}$ represents the label of $p_{u}$. Specifically, it is classified as a positive sample when $s_{u}$ is larger than $t$; otherwise, it is classified as negative. To bring positive samples closer to $a$ and push negative samples farther away, we employ InfoNCE [oord2018representation] to measure the distance between samples:
@@ III-C Synchronized Alignment Fusion (SAF)
Although bimodal features are aligned at deep layers, they still remain slightly offset at high-resolution shallow layers. In addition, the distribution of features varies significantly between modalities due to their distinct physical properties. Effectively fusing these complementary features is essential for achieving BSOD [10315195, 9926193, 10003255, cong2023point, 10042233, 10127616]. To address these challenges, we propose SAF, which uses an FFT-based multiple-filtering strategy to fuse bimodal complementary features and align them in both spatial and channel dimensions.
As shown in Fig. 2, SAF consists of normalization layers, residual connections, a feed-forward network, and a token mixer. To capture the local relevance, we equip a $DConv_{3}(\cdot)$ in the feed-forward network [10015667]. For the token mixer, we introduce the efficient group filter (EGF). In EGF, we first integrate the normalized bimodal features $\tilde{f_{i}^{r}}$ and $\tilde{f_{i}^{t}}$ into the same feature $f_{i}^{a}$, reducing the computational cost for feature alignment and fusion. This process is denoted as:
@@
where ++ is element-wise addition. After applying a StarReLU [yu2022metaformer] layer $SR(\cdot)$, we perform an FFT on the integrated feature $f_{i}^{a}$, then multiply it by the group filter $\mathcal{G}(\cdot)$, where each channel is linearly coupled. The group channel weights in $\mathcal{G}(\cdot)$ are determined by applying global average pooling $GAP(\cdot)$ to $f_{i}^{a}$. The use of channel-weighted $\mathcal{G}(\cdot)$ enhances the sensitivity of capturing salient details and spatial shifts in the spatial dimension, thereby facilitating the fusion and alignment of bimodal features. Finally, the output $\tilde{f_{i}^{s}}$ of EGF is obtained by inverse FFT (IFFT). The whole process is expressed as:
@@
where $\times$ is element-wise multiplication. The $N$ global filters [rao2021global] that make up $\mathcal{G}(\cdot)$ are denoted as $\{G_{1},G_{2},\ldots,G_{N}\}$. The channel-weighted $\mathcal{G}(\cdot)\in\mathbb{C}^{B\times C\times H\times(\frac{W}{2}+1)}$ is defined as:
@@ III-D Decoder
In encoder-decoder structured networks, the distribution of features varies across levels. Low-level features capture basic patterns, such as edges and textures, while high-level features represent more abstract concepts and semantic information [gao2021unified, wang2021cgfnet]. To utilize these diverse representations, we introduce a simple decoder that fits these features gradually from high to low levels. Each decoder layer consists of a $DConv_{3}(\cdot)$, a $Norm(\cdot)$, and ReLU function, denoted as $DNR(\cdot)$. The whole process is represented as:
@@
For the loss function, we employ the commonly used BCE [de2005tutorial] and IoU [mattyus2017deeproadmapper] losses to supervise the learning of $\mathcal{S}$, and employ SCAL to supervise the alignment of $f_{4}^{r}$ and $f_{4}^{t}$, which can be mathematically described as:
@@ IV-A Datasets
To comprehensively validate our proposed AlignSal, we first conducted experiments on the AAV-based unaligned bimodal salient object detection (BSOD) task. To further assess its generalization, we extended our evaluation to three additional bimodal tasks: weakly aligned BSOD, aligned BSOD, and remote sensing change detection (RSCD).
1) AAV-based Unaligned BSOD: We conducted extensive experiments on the AAV RGB-T 2400 dataset [10315195], which contains 2400 instantly captured unaligned RGB-T image pairs falling at different times of the day and different seasons of the year. Each image pair has a corresponding pixel-wise annotation. Following [10315195], all BSOD models were trained on the AAV RGB-T 2400 training set and tested on the testing set.
2) Weakly Aligned BSOD: We employed unaligned-VT821, unaligned-VT1000, and unaligned-VT5000 datasets [tu2022weakly] for the experiments. These three datasets contain 6821 RGB-T image pairs and were obtained by performing random affine transformations from the three corresponding aligned datasets VT821 [wang2018rgb], VT1000 [tu2019rgb], VT5000 [9767629]. Following [10315195, tu2022weakly], all BSOD models were trained on the unaligned-VT5000 training set and tested on the remaining datasets.
3) Aligned BSOD: We utilized VT821 [wang2018rgb], VT1000 [tu2019rgb], and VT5000 [9767629] datasets, which collectively contain 6821 manually aligned RGB-T image pairs in the natural scene. Following [10003255, 10127616], all BSOD models were trained on the VT5000 training set and tested on the remaining datasets.
4) RSCD: To assess the generalization of AlignSal in high-altitude scenarios, we conducted experiments on the LEVIR-CD+ dataset [chen2020spatial], which comprises 985 remote sensing image pairs documenting significant building changes. Each image has an ultra-high resolution of 1024 $\times$ 1024 pixels at 0.5 meters per pixel. All RSCD models were trained on the dataset’s training set and evaluated on its testing set.
@@ IV-B Evaluation Metrics
To fairly evaluate model performance for BSOD tasks, we used six widely recognized evaluation metrics. E-measure [ijcai2018p97] ($Em$) assesses the joint contribution of pixel-level accuracy and image-level matching. S-measure [fan2017structure] ($Sm$) combines region-aware and object-aware evaluation to quantify the structural similarity. F-measure [5206596] ($F_{\beta}$}) balances precision and recall to provide a harmonic mean assessment of accuracy. While weighted F-measure [6909433] ($wF_{\beta}$}) adjusts precision and recall by pixel importance for a refined saliency evaluation. Mean absolute error [6247743] ($\mathcal{M}$) calculates the average pixel-wise absolute difference between prediction and ground truth. We also used Precision-recall (PR) curve to show the trade-off between precision and recall across various thresholds.
When evaluating model performance for the RSCD task, we employed five additional metrics [chen2024changemamba], including recall (Rec.), precision (Pre.), overall accuracy (OA), F1 score (F1), and IoU [mattyus2017deeproadmapper].
@@
TABLE I: Quantitative Comparison of E-measure (EmEm), S-measure (SmSm), F-measure ($F_{\beta}$}), Weighted F-measure ($wF_{\beta}$}), and Mean Absolute Error ($\mathcal{M}$) Between Our AlignSal and Sixteen State-of-the-Art BSOD Models on the AAV RGB-T 2400 Dataset [10315195]. Type A, U, and $U$ Stand for Aligned, Unaligned, and AAV-based Unaligned BSOD Models, Respectively. The Best Results are Labeled Red and the Second Best Results are Labeled Blue. We Tested the FPS of Different Models on an NVIDIA RTX4060 Laptop GPU.
@@
CATNet23 [sun2023catnet] A Swin-B 0.7420 0.8155 0.6331 0.5189 0.0182 262.6M 341.8G 19.7
PICRNet23 [cong2023point] A Swin-T 0.8609 0.7685 0.6129 0.6083 0.0191 112.0M 27.1G 30.3
@@
LAFB24 [wang2024learning] A Res2Net-50 0.9633 0.9040 0.8471 0.8161 0.0068 453.0M 139.7G 29.3
DCNet22 [tu2022weakly] U VGG-16 0.9276 0.8292 0.7500 0.7997 0.0090 24.1M 207.2G 14.5
MROS23 [10315195] $U$ Res2Net-50 0.9747 0.8963 0.8506 0.8494 0.0063 90.4M 45.3G 12.2
AlignSal (Ours) $U$ CDFFormer-S18 0.9654 0.9085 0.8599 0.8277 0.0062 27.1M 22.9G 30.8
@@
$Em\uparrow$ $Sm\uparrow$ $wF_{\beta}\uparrow$ $F_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $\uparrow$ $Em\uparrow$ $Sm\uparrow$ $wF_{\beta}\uparrow$ $F_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $Em\uparrow$ $Sm\uparrow$ $wF_{\beta}\uparrow$ $F_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $Em\uparrow$ $Sm\uparrow$ $wF_{\beta}\uparrow$ $F_{\beta}\uparrow$ $\mathcal{M}\downarrow$
@@
CATNet23 [sun2023catnet] 0.7734 0.8331 0.6627 0.5645 0.0183 0.7511 0.8227 0.6520 0.5393 0.0166 0.7281 0.7953 0.6079 0.5021 0.0207 0.8031 0.8946 0.7514 0.5975 0.0054
PICRNet23 [cong2023point] 0.8678 0.7632 0.6067 0.6233 0.0239 0.8600 0.7671 0.6175 0.6101 0.0185 0.8469 0.7461 0.5842 0.5830 0.0219 0.9206 0.8566 0.7510 0.7240 0.0061
@@
LAFB24 [wang2024learning] 0.9769 0.9142 0.8759 0.8531 0.0057 0.9673 0.9115 0.8650 0.8332 0.0067 0.9576 0.8914 0.8297 0.7990 0.0081 0.9819 0.9511 0.9199 0.8842 0.0020
DCNet22 [tu2022weakly] 0.9529 0.8369 0.7748 0.8229 0.0090 0.9424 0.8314 0.7639 0.8134 0.0089 0.9185 0.8120 0.7285 0.7839 0.0105 0.9740 0.8886 0.8490 0.8813 0.0038
@@ IV-C Implementation Details
We trained our model on an NVIDIA Tesla P100 GPU, using the Adam optimizer [kingma2014adam] with a batch size of 8. Our model utilized the CDFFormer-S18 [tatsunami2024fft] as the encoder, which is pre-trained on ImageNet. The initial learning rate was set to $3\times 10^{-5}$ and decreased by a factor of 10 every 100 epochs, over a total of 300 epochs. The input resolution was set to $384\times 384$.
@@ IV-D Model Comparison
We compared our AlignSal with fourteen state-of-the-art aligned BSOD models, one state-of-the-art unaligned BSOD model, and one state-of-the-art AAV-based unaligned BSOD model. The aligned BSOD models include MIDD [9454273], CSRNet [9505635], MoADNet [jin2022moadnet], SwinNet [9611276], OSRNet [9803225], TNet [9926193], MGAI [10003255], CATNet [sun2023catnet], PICRNet [cong2023point], LSNet [10042233], HRTNet [9869666], CAVER [10015667], WaveNet [10127616], and LAFB [wang2024learning]. The unaligned BSOD model is DCNet [tu2022weakly]. The AAV-based unaligned BSOD model is MROS [10315195]. All saliency maps were obtained from the authors or generated by the open-source code and were evaluated by the Saliency Evaluation Toolbox [^2] https://github.com/jiwei0921/Saliency-Evaluation-Toolbox.
Quantitative Evaluation. To fairly evaluate the performance of our AlignSal on the AAV RGB-T 2400 dataset [10315195], we showed the comparison of $Em$, SmSm, $wF_{\beta}$, $F_{\beta}$, $\mathcal{M}$,
Figure 3: Visual comparison of our AlignSal with state-of-the-art BSOD models in several challenges, including fast AAV movement ($1^{st}$, $2^{nd}$, and $6^{th}$ rows) and fast object movement ($8^{th}$ row) in object blurring scenes, low illumination ($1^{st}$ and $8^{th}$ rows), street light exposure ($3^{rd}$ and $9^{th}$ rows), and extreme low illumination ($4^{th}$ row) in illumination changing scenes, small objects ($2^{nd}$, $4^{th}$, $5^{th}$, and $8^{th}$ rows), out-of-view ($1^{st}$, $2^{nd}$, $6^{th}$, and $8^{th}$ rows), multiple objects($1^{st}$-$4^{th}$, $7^{th}$-$9^{th}$ rows), scale variation ($8^{th}$ row), and center bias ($4^{th}$ and $6^{th}$ rows) in object changing scenes, rain ($8^{th}$ row) and snow ($9^{th}$ row) in weather changing scenes. GT represents ground truth.
Figure 4: (a) Precision-recall (PR) and (b) F-measure-threshold (FT) curves of BSOD models on the AAV RGB-T 2400 dataset.
and three complexity metrics (i.e., Params, FLOPs, and FPS) of our AlignSal with the above BSOD models in Table I. Notably, with real-time inference speed and lower complexity, our model achieved superior performance compared to the state-of-the-art BSOD models. Specifically, compared to the existing top-performing AAV-based unaligned model MROS [10315195], our model surpassed it in three out of five metrics (i.e. an improvement of 1.4%, 1.1%, and 1.6% in $Sm$, $wF_{\beta}$, and $\mathcal{M}$, respectively), while being much lighter (i.e. 70% reduction in Params) and having much faster inference speed (i.e. half the FLOPs and 2.5 times the FPS). Compared to the existing best aligned BSOD model LAFB [wang2024learning], our model outperformed it by 0.2%, 0.5%, 1.5%, 1.4%, and 8.8% in $Em$, SmSm, $wF_{\beta}$, $F_{\beta}$, and $\mathcal{M}$, respectively, while having a 94.1% and 83.6% reduction of Params and FLOPs. In addition, we also provided the comparison of PR and F-measure-threshold (FT) curves in Fig. 4. We can observe that the PR curve of our model is closest to (1,1), and the area under the FT curve is the largest as well.
@@
Qualitative Comparison. As shown in Fig. 3, we present a visual comparison of our AlignSal with existing state-of-the-art BSOD models across several challenging scenes, including fast AAV movement ($1^{st}$, $2^{nd}$, and $6^{th}$ rows) and fast object movement ($8^{th}$ row) in object blurring scenes; low illumination ($1^{st}$ and $8^{th}$ rows), street light exposure ($3^{rd}$ and $9^{th}$ rows), and extreme low illumination ($4^{th}$ row) in illumination changing scenes; small objects ($2^{nd}$, $4^{th}$, $5^{th}$, and $8^{th}$ rows), out-of-view ($1^{st}$, $2^{nd}$, $6^{th}$, and $8^{th}$ rows), multiple objects($1^{st}$-$4^{th}$, $7^{th}$-$9^{th}$ rows), scale variation ($8^{th}$ row), and center bias ($4^{th}$ and $6^{th}$ rows) in object changing scenes; rain ($8^{th}$ row) and snow ($9^{th}$ row) in weather changing scenes. As shown, a scene often contains multiple challenges and AlignSal can completely capture objects in these extreme scenes. This is attributed to our proposed SCAL and SAF and the way they cooperate within AlignSal.
TABLE IV: Ablation Studies of the Key Components in AlignSal on the AAV RGB-T 2400 Dataset. The Best Results are Labeled Red.
@@
SCAL aligns and mutually refines bimodal features at the semantic level using a contrastive strategy, while SAF aligns the two modalities at the pixel level by using the powerful filtering mechanism to effectively explore the bimodal complementary information and capture fine details.